# Magneto-inertial range dominated by magnetic helicity in space plasmas

### v3

@@ Abstract
Magneto-inertial range dominated by magnetic helicity has been studied using results of the numerical simulations, laboratory measurements, solar, solar wind, and the Earth’s and planets’ magnetosphere observations (spacecraft measurements). The spectral data have been compared with the theoretical results based on the distributed chaos notion in the frames of the Kolmogorov-Iroshnikov phenomenology. The transition from magnetohydrodynamics to kinetics in the electron and Hall magnetohydrodynamics, and in a fully kinetic 3D approach, as well as in the solar wind, solar photosphere, and at the special events (reconnections, Kelvin-Helmholtz instability, isolated flux tube interchanges, etc.) in the magnetosphere of Earth, Saturn, Jupiter, and Mercury has been also discussed.
@@ I Introduction
In Section X conclusions summarizing the results obtained in the previous sections will be presented.
@@ IV Magneto-inertial range of scales
$B_{c}\propto\varepsilon_{h}^{1/2}~{}\varepsilon^{1/6}~{}k_{c}^{1/6}$ (16)
@@
$B_{c}\propto\varepsilon_{\hat{h}}^{1/2}~{}(\varepsilon\widetilde{B}_{0})^{1/8}k_{c}^{1/8}$ (17)
with the modified magnetic helicity (see the Eq. (2)) dissipation rate $\varepsilon_{\hat{h}}$. However, the dimensional estimation Eq. (16) can be still valid for some cases even with a strong external magnetic field.
@@ IX Solar active regions
X Conclusions
@@
Comparing the spectral behavior in the solar wind for different distances from the Sun one can conclude that in the near Sun heliosphere, the role of the background magnetic field becomes more prominent Eq. (23) than for the larger distances from the Sun Eq. (22). The transitional distance for this phenomenon seems to be near the value 1.4 AU Fig. 20.
In the planets’ magnetosphere, the spectral law Eq. (23) is also prevailing, i.e. the role of the background magnetic field is generally significant there as well (and also in the intensive solar active regions the role of the background magnetic field can become significant at a mature stage of their development).
XI Acknowledgments

### v4

@@ Abstract
Magneto-inertial range dominated by magnetic helicity has been studied using results of numerical simulations, laboratory measurements, solar, solar wind, the Earth’s and planets’ magnetosphere observations (spacecraft measurements), and the global magnetic observatory network. The spectral data have been compared with the theoretical results based on the distributed chaos notion in the frames of the Kolmogorov-Iroshnikov phenomenology. The transition from magnetohydrodynamics to kinetics in the electron and Hall magnetohydrodynamics, and in a fully kinetic 3D approach, as well as in the solar wind, solar photosphere, and at the special events (reconnections, Kelvin-Helmholtz instability, isolated flux tube interchanges, etc.) in the magnetosphere of Earth, Saturn, Jupiter, and Mercury has been studied using the above-mentioned data. Despite the considerable differences in the physical parameters and scales, the results of numerical simulations are in quantitative agreement with the observational data in the frames of the magneto-inertial range notion. Temporal variability of the magnetic field at Earth’s surface under the influence of the ionosphere, magnetosphere, and solar wind has been also briefly discussed in this context.
@@ I Introduction
In a wide range of time scales temporal variability of the magnetic field at Earth’s surface is dominated by the variability of the external magnetic field arising from the interaction of the near-Earth solar wind with the Earth’s magnetosphere and its inner edge - ionosphere. In Section X this variability will be studied using the magneto-inertial range of scales notion and the data obtained with the global magnetic observatory network.
In Section XI conclusions summarizing the results obtained in the previous sections will be presented.
@@ IV Magneto-inertial range of scales
$B_{c}\propto\varepsilon_{h_{m}}^{1/2}~{}\varepsilon^{-1/6}~{}k_{c}^{1/6}$ (16)
@@
$B_{c}\propto\varepsilon_{\hat{h}_{m}}^{1/2}~{}(\varepsilon\widetilde{B}_{0})^{-1/8}k_{c}^{1/8}$ (17)
with the modified magnetic helicity (see the Eq. (2)) dissipation rate $\varepsilon_{\hat{h}_{m}}$. However, the dimensional estimation Eq. (16) can be still valid for some cases even with a strong external magnetic field.
@@ IX Solar active regions
X Temporal variability of the magnetic field at Earth’s surface
Until now the spatial properties of the magneto-inertial range have been studied (the above-discussed frequency spectra reflect these properties due to the Taylor’s ‘frozen-in’ hypothesis). It is interesting also to study the true temporal spectral properties of the magneto-inertial range. Temporal variability of the magnetic field at Earth’s surface provides such a possibility. In a wide range of time scales this variability is dominated by the variability of the external magnetic field arising from the interaction of the near-Earth solar wind with Earth’s magnetosphere and its inner edge - ionosphere (see, for instance, a recent paper cc and references therein).
To find the true frequency (temporal) spectra corresponding to the magneto-inertial range one should replace the Eqs. (16), (17) and (18) by equations
$B_{c}\propto\varepsilon_{h_{m}}^{1/2}~{}\varepsilon^{-1/4}~{}f_{c}^{1/4},$ (32)
$B_{c}\propto\varepsilon_{\hat{h}_{m}}^{1/2}~{}(\varepsilon\widetilde{B}_{0})^{-1/6}f_{c}^{1/6},$ (33)
and
$B_{c}\propto f_{c}^{\alpha}$ (34)
where $f_{c}$ is a true characteristic frequency.
Hence $\alpha=1/4$ for the case represented by Eq. (32) and $\alpha=1/6$ for the case represented by Eq. (33). It follows from the Eq. (21) (which is obviously valid for the frequency spectra as well) that $\beta=1/3$ for the case of the Eq. (32) and $\beta=1/4$ for the case of the Eq. (33). Corresponding frequency spectra have the form
$E(f)\propto\exp-(f/f_{\beta})^{1/3}$ (35)
and
$E(f)\propto\exp-(f/f_{\beta})^{1/4}$ (36)
for the Eq. (32) and Eq. (33) respectively.
Figure 39 shows the true frequency magnetic energy spectrum obtained using measurements provided by the global magnetic observatory network (the spectral data were taken from Fig. 1 of the paper cc ). The range of frequencies shown in the Fig. 39 corresponds to the above-mentioned time scales for which variability of the magnetic field at Earth surface is dominated by the variability of the external magnetic field arising from the interaction of the near-Earth solar wind with Earth’s magnetosphere and ionosphere.
The 27-day peak and its harmonics are related to the solar rotation period, whereas the peak and its harmonics corresponding to daily variations are presumably associated with thermal electric currents in the ionosphere on the heated day side of the globe. Seasonal variations of the ionosphere heating can be the main reason for the appearance of the spectral peaks corresponding to the 1 year and 6 months (see Ref. cc and references therein).
The dashed curve in the Fig. 39 indicates correspondence to the spectral law Eq. (36) taking into account the influence of a mean magnetic field on the magneto-inertial range of scales.
XI Conclusions
@@
Figure 39: The true frequency magnetic energy spectrum obtained using measurements provided by the global observatory network.
Comparing the spectral behavior in the solar wind for different distances from the Sun one can conclude that in the near Sun heliosphere, the role of the background mean magnetic field becomes more prominent Eq. (23) than for the larger distances from the Sun Eq. (22). The transitional distance for this phenomenon seems to be near the value 1.4 AU Fig. 20.
In the planets’ magnetosphere, the spectral law Eq. (23) is also prevailing, i.e. the role of the background mean magnetic field is generally significant there as well (and also in the intensive solar active regions the role of the background mean magnetic field can become significant at a mature stage of their development). The analogous situation takes also place for the temporal variability of the magnetic field at Earth’s surface dominated by the variability of the external magnetic field arising from the interaction of the near-Earth solar wind with Earth’s magnetosphere and ionosphere.
Despite the considerable differences in the physical parameters and scales, the results of numerical simulations are in quantitative agreement with the observational data in the frames of the magneto-inertial range notion.
XII Acknowledgments